OpenStudy (anonymous):

what is 100.5 in pi terms

OpenStudy (anonymous):

\[\frac{ 100.5 }{ \pi } \approx 31.99\] So \[40\pi\] perhaps?

OpenStudy (gorv):

we have toconert it into radians

OpenStudy (gorv):

\[\frac{ \pi }{ 180 }*100.5\]

OpenStudy (gorv):

0.55 pi approx 0.5 pi